OpenStudy (anonymous):

A system of equations is shown below: 2x = 5y + 4 3x - 2y = -16 What is the solution to this system of equations?

OpenStudy (anonymous):

try solving for x in the first equation and then use substitution.

OpenStudy (anonymous):

What do you mean? I'm terrible at this.

OpenStudy (anonymous):

@Redemption

OpenStudy (anonymous):

2x = 5y + 4 solve for the x variable

OpenStudy (anonymous):

How man.... -2 from every side?

OpenStudy (anonymous):

2x = 2 times x to get x by itself you do the opposite which is division so you dvide by 2

OpenStudy (anonymous):

So, x = 5y + 2?

OpenStudy (anonymous):

@Redemption Can you actually help?

OpenStudy (anonymous):

x = 5/2 y + 2 now substitute.

OpenStudy (anonymous):

@Redemption So, x = 2.5y + 2 === X = 4.5?

OpenStudy (anonymous):

With that being the occsion, I'd assume the answer is (4, 8)

OpenStudy (jdoe0001):

hmm @A.Squad you're supposed to know how to do linear simplification... thus

OpenStudy (anonymous):

I understand it and all I'm just terrible at it.

OpenStudy (jdoe0001):

well... still.. the exercise is just a system of equations.... if you dunno linear simplification, it wouldn't help give you this exercise

OpenStudy (anonymous):

Okay, so I should just do.. 2x = 5y + 4 x = 2.5 + 2 X = 4.5?

OpenStudy (anonymous):

What did you do with the y - value?

OpenStudy (anonymous):

How it's 2x = 5+.... OHOHHHH

OpenStudy (anonymous):

2X = 5Y + 4 2/2x = 5y + 4 x = 5y + 4?

OpenStudy (anonymous):

I mean x = 5y +2?

OpenStudy (jdoe0001):

well... more or less, yes but you may want to keep it in fraction form usually so \(\bf 2{\color{brown}{ x}} = 5y + 4\\ 3{\color{brown}{ x}} - 2y = -16 \\ \quad \\ \quad \\ 2{\color{brown}{ x}} = 5y + 4\implies {\color{brown}{ x}}=\cfrac{5y + 4}{2}\implies {\color{brown}{ x}}=\cfrac{5y }{2}+\cfrac{\cancel{ 4 }}{\cancel{ 2 }} \\ \quad \\ {\color{brown}{ x}}=\cfrac{5y }{2}+2\qquad thus \\ \quad \\ 3{\color{brown}{ x}} - 2y = -16\implies 3{\color{brown}{ \cfrac{5y }{2}+2}} - 2y = -16\) notice.... once we know what "x" is then you can just substitute on the 2nd equation for it, and solve for "y"

OpenStudy (jdoe0001):

and once you know what "y" is, as a constant then you can plug that in on either equation, to get "x" as a constant

OpenStudy (anonymous):

Okay, hold up, lemme studdy this for a sec.

OpenStudy (anonymous):

Ohhh.

OpenStudy (anonymous):

Wait, now I'm just confused... Which would it be? (-8, -4) (8, 4) (-4, -8) (4, 8)

OpenStudy (jdoe0001):

heheh nevermind the choices.... what matters is the process the choices are there as guidelines, no authority supposed the choices are wrong say how much is 2+2 choices -> 25, 3, 7, 11 so...the choices shouldn't matter, the choices MUST match your solution, not the other way around

OpenStudy (anonymous):

Well, those are the answers.

OpenStudy (jdoe0001):

tis ok.... so... simplify the substitution above to get "y" :)

OpenStudy (anonymous):

Alright lemme do this.

OpenStudy (anonymous):

So, x is 4?

OpenStudy (radar):